OpenStudy (anonymous):

A board 60 in. long is cut into two parts so that the longer piece is 5 times the shorter. What are the lengths of the two pieces of board?

OpenStudy (anonymous):

L+S=60 L=5S 5S+S=60 6S=60 S=10 L=5*10=50